If (2x – 2)(2 – x) = 0, what are all the possible values of x?
A. 0 only B. 1 only C. 2 only D. 1 and 2 only E. 0, 1, and 2

Answer by jim_thompson5910(35256)

If (2x – 2)(2 – x) = 0, then 2x-2=0 or 2-x=0. Solve for x in each equation to get x=1 or x=2


So the answer is D)


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim